Assume that shyness, measured on a 1 to 10 scale, is normally distributed in a population. A statistician would guess that the shyness level of 10 randomly selected people is:

A. one standard deviation above the mean

B. the population mean minus n.

C. the mean of the population.

D. one standard deviation below the mean

If r = –.69, SX = 1.08, and SY = 6.98, what is the slope?

A. -0.11

B. 0.11

C. -4.46

D. 10.93

A vocational counselor uses the number of days without employment to predict her clients' feelings of self efficacy, measured on a scale of 1 to 5, with higher numbers meaning that clients feel more secure in their job related abilities. The slope of the regression line is –1.02. Which statement is the best interpretation for this finding?

A. The decrease in self-efficacy of 1.02 points is caused by each additional day of unemployment.

B. The least number of days a person can be unemployed and still feel self-efficacious is 3.98 points

C. For every 1-point increase in self-efficacy, there is an associated decrease in the number of days of unemployment.

D. For every additional day of unemployment, there is an associated decrease in self-efficacy of 1.02 points.

If the Y-intercept is negative, the best-fitting line passes through the Y-axis:

A. above zero

B. close to zero

C. below zero

D. far from zero
Answer #1

(1)

Question:

Assume that shyness, measured on a 1 to 10 scale, is normally distributed in a population. A statistician would guess that the shyness level of 10 randomly selected people is:

Correct option:

C.    the mean of the population

Explanation:
By Central Limit Thoerem, the sampling distribution of sample means is Normal Distribution with mean = mean of the population.

(2)

Question:

If r = –.69, SX = 1.08, and SY = 6.98, what is the slope?

Correct option:

C.   - 4.46

Explanation:

Slope (b) is given by:

b=r\times \frac{S_{Y}}{S_{X}}=-0.69\times \frac{6.98}{1.08}=-4.46

(3)

Question:

A vocational counselor uses the number of days without employment to predict her clients' feelings of self efficacy, measured on a scale of 1 to 5, with higher numbers meaning that clients feel more secure in their job related abilities. The slope of the regression line is –1.02. Which statement is the best interpretation for this finding?

Correct option:

A     The decrease in self-efficacy of 1.02 points is caused by each additional day of unemployment.

(4)

Question:

If the Y-intercept is negative, the best-fitting line passes through the Y-axis:

Correct option:

C.    below zero